# Physics beyond the Standard Model with the DSA-2000

### v1

@@ 2.4 Axion clouds around neutron stars
It was recently shown that NSs are efficient emitters of axions in the radio frequency range [124]. The mechanism for axion generation is related to the formation of vacuum gaps in NS magnetospheres. Vacuum gaps host large, accelerating electric fields $E_{\parallel}$, parallel to the background magnetic field, and are believed to be the sites of non-thermal electromagnetic emission. A large $E_{\parallel}$ is unstable to runaway $e^{\pm}$ pair creation, which dynamically screens $E_{\parallel}$, giving rise to a space and time-varying ${\bf E}\cdot{\bf B}$, which acts as a classical source of axions, according to the Klein-Gordon equation [124],
@@
Dynamical screening of $E_{\parallel}$ in pulsar polar caps is believed to be the origin of pulsed radio emission observed in rotating neutron stars [125]. Axions are produced with a broad spectrum, set by the spectrum of plasma oscillations in the gap. A fraction of produced axions resonantly convert to photons, giving rise to broadband radio emission. Non-observation of these radio signals from NSs has yielded some of the most stringent constraints on axions with masses between $10^{-8}\ {\rm eV}\lesssim m_{a}\lesssim 10^{-5}$ eV [106]. If the axion mass lies in the range $10^{-9}$ eV $\lesssim m_{a}\lesssim 10^{-4}$ eV, a significant fraction of produced axions remain gravitationally bound to the NS [28]. The population of bound axions forms a cloud whose density can grow over astrophysical timescales, reaching densities over twenty orders of magnitude larger than the local DM density ($\rho_{\rm DM}\approx 0.4$ GeV/cm3[28]. The cloud also experiences dissipation through a number of processes, with the dominant effect being resonant axion-to-photon conversion at large distances. The dissipation of the cloud gives rise to bright, narrow-band radio emission that may be observable with sensitive telescopes and arrays such as the DSA-2000. The radio lines emanating from axion clouds have a fractional bandwidth of $\mathcal{B}/f\sim\mathcal{O}(10^{-2})$, which is set by the kinematic requirement for resonant axion-to-photon conversion as described in detail below. This signal is much broader than the expected signal associated with DM axion conversion in the magnetosphere. Crucially, the line-width of radio emission from axion clouds exceeds the broadening from individual NS motions, thus allowing a simple search strategy from populations of NSs. Below we estimate the sensitivity of DSA-2000 to radio signals from axion clouds.
In this subsection, we review the semi-analytic formalism for axion production in pulsar polar caps, with a focus on the gravitationally bound population [28, 29]. We then derive analytic estimates for the luminosity and bandwidth of the associated radio signal arising from axion cloud dissipation. Finally, we use these estimates to forecast the sensitivity of DSA-2000 to both individual pulsars and the broader population expected to be discovered by the array. A visual summary of the formation of axion clouds and their associated radio signals is shown in Fig. 5.
@@
The axions produced by dynamical screening of the accelerating electric field can be obtained by solving the axion equation of motion (2.15), giving a differential production rate of [126]:
@@
where $k\cdot x=-\omega_{k}t+\vec{k}\cdot\vec{x}$, and $\omega_{k}^{2}=\vec{k}^{2}+m_{a}^{2}$. The time integral in (2.17) is performed over the limit cycle of the pair cascade, and the spatial integral is performed over the volume of the polar cap, here approximated as a cylinder of radius $r_{\rm pc}=R_{\rm NS}\sqrt{\Omega R_{\rm NS}}$, where $\Omega=2\pi/P$ is the rotation frequency of the pulsar, and height [127]
@@
where $\omega_{k}=\sqrt{k^{2}+m_{a}^{2}}$, $\tilde{x}=x/r_{\rm pc}$, $\tilde{y}=y/r_{\rm pc}$, $\tilde{z}=z/h$, $\tilde{t}=t/t_{\rm open}$, $t_{\rm open}$ is the typical timescale over which the gap is open, and $({\bf E}\cdot{\bf B})_{0}=2\Omega hB_{0}^{2}$ is the maximum value of ${\bf E}\cdot{\bf B}$ in the gap. The gap opening timescale is related to the light crossing time of the gap, $t_{\rm open}=h$. The duty cycle of the gap opening, $t_{\rm open}/T$ is taken to be $10\%$ as in [29]. Equation (2.19) can be rewritten as
@@
The function $\tilde{g}(\mathbf{k})$ depends on both the spectrum of spatial fluctuations in ${\bf E}\cdot{\bf B}$ and the axion mass. The former has two major components: the first coming from the geometry of the gap, and the second from small-scale plasma oscillations that arise due to non-stationary pair creation and dynamical screening in the gap [29]. For light axions, with wavelengths larger than the characteristic gap size, $h$, the geometric contribution to $\tilde{g}(\mathbf{k})$ dominates. In this low-momentum limit, $|\tilde{g}(\mathbf{k})|\to 1$. For heavy axions, with wavelengths much shorter than the typical gap height, axion production depends on details of the plasma oscillation spectrum, understanding which requires kinetic particle-in-cell (PIC) simulations. Semi-analytic modeling of the plasma structure during the discharge phase shows a frequency spectrum $E_{\parallel}(\omega)\approx E_{\parallel,0}(\omega/\omega_{0})^{-1/2}$, where $\omega_{0}$ is an IR scale, taken to be $\omega_{0}\sim 1/h$. Assuming spatial fluctuations have the same spectrum [^5] See [29] for a plausibility argument for this assumption and a statement of the requirement of simulations for definitive justification.[29] demonstrated that in the small-$k$ limit, relevant for gravitationally bound axions, the effect of small-scale oscillations leads to a suppression of $\sim(m_{a}h)^{-4}$ in the axion luminosity [^6] The expression quoted in [29] is $(m_{a}h)^{-3}(\omega_{0}/m_{a})$, but we assume $\omega_{0}\sim 1/h$.. Putting together the high-mass and low-mass regimes, we obtain
@@
where $\Theta(x)$ is the Heaviside theta function, which selects axions with velocity less than the escape velocity of the neutron star, $v_{\rm esc}=\sqrt{2GM_{\rm NS}/R_{\rm NS}}\approx 0.5c$. A simplifying assumption made in (2.22) is that the differential axion production rate is approximately flat at small $k$. The distribution of axion velocities near the NS surface was studied using ray-tracing simulations in [29], and is approximately flat at low-$k$, before dropping at about $v\approx 0.9v_{\rm esc}$. Such simulations are required to make more robust predictions about axion cloud formation and dissipation. The total luminosity of axions emitted into bound states is then
@@
Equation (2.23) agrees with the result in [29] up to an order unity factor that arises from the approximation $d^{3}k\sim 4\pi\bar{k}^{3}/3\sim 4\pi m_{a}^{3}/81$, where $\bar{k}\sim m_{a}/3$ is the average cloud axion momentum.
@@
The axion cloud signal appears as a broadened spectral line centered at frequency corresponding to the axion mass, with bandwidth set by the axion velocity distribution (see Eq. (2.26)). For generality, the scaling equation above has assumed a constant signal bandwidth $\mathcal{B}\sim 60\,\text{MHz}$, suitable for both a population survey and a fixed target observation, the latter case could be made more precise by incorporating the full parametric dependence as given in Eq. (2.26). Note that in the DSA-2000 frequency range $0.7\,\mathrm{GHz}\lesssim f\lesssim 2\,\mathrm{GHz}$ (corresponding axion masses $2.9\,\mu\mathrm{eV}\lesssim m_{a}\lesssim 8.3\,\mu\mathrm{eV}$), the axion wavelength is almost always less than the gap height $h$ estimated by Eq. (2.18), such that taking $\min{(1,{(m_{a}h)}^{-4})}\approx{(m_{a}h)}^{-4}$ is justified in the estimate of $\dot{E}_{a}$  [^7] In this regime, an additional radio line is sourced due to the oscillating current associated with the axion field, which may have interesting observational consequences [29]. The magnitude of the signal is primarily determined by the axion-photon coupling strength, the NS surface magnetic field, its distance to Earth, and its rotation period. Optimal candidates are nearby, fast-spinning pulsars with strong magnetic fields.
@@
The preceding section focuses on radio signals from individual NSs. Another approach to axion detection involves searching for signals from populations of NSs [7, 17, 128]. An advantage of population signals is that the total received flux is bounded from below by the flux from a single NS. However, in the case of axion DM, the signal is Doppler broadened to $\mathcal{B}/f\sim v_{\rm gal}/c\approx 7\times 10^{-4}$, where $v_{\rm gal}$ is the galactic velocity dispersion. Therefore, the expected signal is much broader than that of an individual NS’s $\mathcal{B}/f\sim v_{0}\approx 10^{-6}$ [^8] Even signals from individual NSs are expected to have bandwidth greater than $\sim 10^{-6}$ due to the NS’s rotation (see, e.g., [13]), but still less than the Doppler-broadened population signal.. Here, we consider radio signals from a population of NSs surrounded by axion clouds. The expected bandwidth for axion cloud signals in (2.26) dominates over Doppler broadening. Therefore, we can estimate the total flux density received from a population of NSs by simply summing the luminosity density of each individual NS:
@@
Figure 6: Distribution plots of the population model (dots), the ATNF catalog (triangles), the Crab pulsar (circled triangle), and the best predicted pulsar to be discovered by the DSA-2000 (circled dot). The color bar shows the axion cloud signal flux density from each individual source, normalized to the flux density from the Crab pulsar. The grey lines show the contours of the axion cloud signal flux density in the corresponding parameter space, normalized to the axion cloud signal flux density from the Crab pulsar. The contour lines are calculated with the varying parameters shown in the plots ($B_{0},d$ for the left panel and $B_{0},P$ for the right panel) and the rest of the parameters fixed to be identical to that of the Crab pulsar. We have taken benchmark values of $m_{a}=5\,\mu\text{eV}$ and $g_{a\gamma\gamma}=10^{-15}\,\text{GeV}^{-1}$ in our estimation of the flux, and the Crab pulsar signal flux density is $F_{\text{Crab}}=1.6\times 10^{-9}\,\text{Jy}$. The left panel shows the distribution in surface magnetic field ($B_{0}$ [G]) and distance to Earth ($d$ [pc]). The distribution plot is bounded from above by the size of the Milky Way, at $d\sim 1.5\times 10^{4}$ pc. The right panel shows the distribution in surface magnetic field ($B_{0}$ [G]) and rotation period ($P$ [s]). The gray region is where the conversion radius $r_{c}$ falls inside the NS. We have discarded NSs older than $10^{7}$ yrs in the population model due to their deviations in the magnetosphere from the GJ model. This age cut keeps the population (dots) away from the region with high period $P$ and weak magnetic field $B_{0}$.
@@
For our population modeling, we rely on a pulsar population snapshot generated from the PrsPopPy simulation [129], which provides NS parameters, including their Period $P$ and distance $d$, and only includes objects discoverable with DSA-2000. We generate their present-day surface magnetic field, $B_{0}\,$[G], by drawing from a log-normal distribution centered at $11.98$ and with a standard deviation of $0.52$. These parameters are consistent with the PsrPopPy simulation [129], as well as the ATNF catalog [71]. Since the magnetosphere of old pulsars is not well-described by the GJ model [130], we discard the NSs with age $t_{\text{NS}}\geq 10^{7}\,\text{yrs}$. We also do not include millisecond pulsars with rotation periods $P\lesssim 0.05\,\text{s}$. After imposing the above additional criteria on the population model for new detections, we are left with $\approx 9000$ pulsars that satisfy those criteria, which we use for the axion cloud forecast. In Fig. 6, we show the distributions of the expected detection population (dot) along with the ATNF catalog (triangle). The axion cloud signal flux density from these sources, normalized to the Crab pulsar (circled triangle), is shown in the color bar. The left panel shows the distribution in surface magnetic field, $B_{0}$ and distance $d$, where we have cut off NSs outside of the Milky Way ($d\gtrsim 1.5\times 10^{4}\,$pc). The right panel shows the distribution in surface magnetic field and pulse period $P$, where we cut off the region where the conversion surface falls inside the NS surface ($r_{c}<R_{\mathrm{NS}}$), as well as NSs older than $10^{7}\,$yrs. We also plot contours of the signal flux density $F_{\mathrm{NS}}$ in the corresponding parameter space with benchmark parameters $m_{a}=5\,\mu\mathrm{eV}$ and $g_{a\gamma\gamma}=10^{-15}\,\mathrm{GeV}^{-1}$ (while fixing the other parameters to be identical to the Crab), normalized against the Crab pulsar ($F_{\mathrm{Crab}}\approx 1.6\times 10^{-9}\,\mathrm{Jy}$). We have chosen the above described cuts on distance, period, and magnetic field for the NSs included in our forecast to be conservative. However, once DSA-2000 discovers new pulsars, the assumptions on the new population made here can be updated and validated before using the collected data to set a constraint.
@@
We also show the best predicted target from the population model (circled dot, parameters shown in Table 3), which represents a prospective candidate that could be discovered by the DSA-2000, whose axion cloud signal flux density is higher than that of the Crab pulsar but has not been observed before. The best predicted target from the population model has a stronger magnetic field, which enables a stronger axion cloud signal than the Crab pulsar. However, due to its larger distance from the Earth, the new candidate produces a smaller non-axion radio flux density than the Crab pulsar [131]. We expect data from the DSA-2000 to discover targets farther from the Earth that could be potential better candidates for observing axion signals. We note that the parameters shown in Table 3 for the best predicted target, informed by our population modeling, are speculative. The data to be taken with the DSA-2000 will ultimately unveil the best discovered target, whose measured parameters will determine the sensitivity to the QCD axion band that can be reached with targeted follow-up observations.
@@
Figure 7: Projected QCD axion discovery reach (SNR $\geq 5$) with the DSA-2000 from radio emission of axion clouds around NSs. We make projections of the Crab pulsar (dashed magenta, $\Delta t_{\text{obs}}=10$ hours), the ATNF population (solid green), and the population model forecasting the DSA-2000 detection (solid blue), based on the planned all sky survey which will observe each NS for 4 hours. We also include projections of a predicted candidate that could potentially be discovered by the DSA-2000 (see Sec. 2.4 for more discussions), with a $\Delta t_{\text{obs}}=10$ hours (dashed red) and a $\Delta t_{\text{obs}}=1000$ hours (dotted red) observation. Other relevant astrophysical constrains are plotted in darker gray (dominated by unbounded polar cap axions from NS [106]), and relevant haloscope constraints (ADMX [80, 81, 82, 83, 84, 85, 86], CAPP [91, 92, 93, 94, 95, 96, 97, 98, 99, 100], RBF+UF [87, 88, 89, 90]) on axion DM are plotted in lighter gray. The QCD axion parameter space is shaded in yellow [111].

### v2

@@ 2.4 Axion clouds around neutron stars
It was recently shown that NSs are efficient emitters of axions in the radio frequency range [Prabhu:2021zve]. The mechanism for axion generation is related to the formation of vacuum gaps in NS magnetospheres. Vacuum gaps host large, accelerating electric fields $E_{\parallel}$, parallel to the background magnetic field, and are believed to be the sites of non-thermal electromagnetic emission. A large $E_{\parallel}$ is unstable to runaway $e^{\pm}$ pair creation, which dynamically screens $E_{\parallel}$, giving rise to a space and time-varying ${\bf E}\cdot{\bf B}$, which acts as a classical source of axions, according to the Klein-Gordon equation [Prabhu:2021zve],
@@
Dynamical screening of $E_{\parallel}$ in pulsar polar caps is believed to be the origin of pulsed radio emission observed in rotating neutron stars [Philippov2020]. Axions are produced with a broad spectrum, set by the spectrum of plasma oscillations in the gap. A fraction of produced axions resonantly convert to photons, giving rise to broadband radio emission. Non-observation of these radio signals from NSs has yielded some of the most stringent constraints on axions with masses between $10^{-8}\ {\rm eV}\lesssim m_{a}\lesssim 10^{-5}$ eV [Noordhuis:2022ljw]. If the axion mass lies in the range $10^{-9}$ eV $\lesssim m_{a}\lesssim 10^{-4}$ eV, a significant fraction of produced axions remain gravitationally bound to the NS [Noordhuis:2023wid]. The population of bound axions forms a cloud whose density can grow over astrophysical timescales, reaching densities over twenty orders of magnitude larger than the local DM density ($\rho_{\rm DM}\approx 0.4$ GeV/cm3[Noordhuis:2023wid]. The cloud also experiences dissipation through a number of processes, with the dominant effect being resonant axion-to-photon conversion at large distances. The dissipation of the cloud gives rise to bright, narrow-band radio emission that may be observable with sensitive telescopes and arrays such as the DSA-2000. The radio lines emanating from axion clouds have a fractional bandwidth of $\mathcal{B}/f\sim\mathcal{O}(10^{-2})$, which is set by the kinematic requirement for resonant axion-to-photon conversion as described in detail below. This signal is much broader than the expected signal associated with DM axion conversion in the magnetosphere. Crucially, the line-width of radio emission from axion clouds exceeds the broadening from individual NS motions, thus allowing a simple search strategy from populations of NSs. Below we estimate the sensitivity of DSA-2000 to radio signals from axion clouds.
In this subsection, we review the semi-analytic formalism for axion production in pulsar polar caps, with a focus on the gravitationally bound population [Noordhuis:2023wid, Caputo:2023cpv]. We then derive analytic estimates for the luminosity and bandwidth of the associated radio signal arising from axion cloud dissipation. Finally, we use these estimates to forecast the sensitivity of DSA-2000 to both individual pulsars and the broader population expected to be discovered by the array. A visual summary of the formation of axion clouds and their associated radio signals is shown in Fig. 5.
@@
The axions produced by dynamical screening of the accelerating electric field can be obtained by solving the axion equation of motion (2.15), giving a differential production rate of [Peskin:1995ev]:
@@
where $k\cdot x=-\omega_{k}t+\vec{k}\cdot\vec{x}$, and $\omega_{k}^{2}=\vec{k}^{2}+m_{a}^{2}$. The time integral in (2.17) is performed over the limit cycle of the pair cascade, and the spatial integral is performed over the volume of the polar cap, here approximated as a cylinder of radius $r_{\rm pc}=R_{\rm NS}\sqrt{\Omega R_{\rm NS}}$, where $\Omega=2\pi/P$ is the rotation frequency of the pulsar, and height [RudermanSutherland]
@@
where $\omega_{k}=\sqrt{k^{2}+m_{a}^{2}}$, $\tilde{x}=x/r_{\rm pc}$, $\tilde{y}=y/r_{\rm pc}$, $\tilde{z}=z/h$, $\tilde{t}=t/t_{\rm open}$, $t_{\rm open}$ is the typical timescale over which the gap is open, and $({\bf E}\cdot{\bf B})_{0}=2\Omega hB_{0}^{2}$ is the maximum value of ${\bf E}\cdot{\bf B}$ in the gap. The gap opening timescale is related to the light crossing time of the gap, $t_{\rm open}=h$. The duty cycle of the gap opening, $t_{\rm open}/T$ is taken to be $10\%$ as in [Caputo:2023cpv]. Equation (2.19) can be rewritten as
@@
The function $\tilde{g}(\mathbf{k})$ depends on both the spectrum of spatial fluctuations in ${\bf E}\cdot{\bf B}$ and the axion mass. The former has two major components: the first coming from the geometry of the gap, and the second from small-scale plasma oscillations that arise due to non-stationary pair creation and dynamical screening in the gap [Caputo:2023cpv]. For light axions, with wavelengths larger than the characteristic gap size, $h$, the geometric contribution to $\tilde{g}(\mathbf{k})$ dominates. In this low-momentum limit, $|\tilde{g}(\mathbf{k})|\to 1$. For heavy axions, with wavelengths much shorter than the typical gap height, axion production depends on details of the plasma oscillation spectrum, understanding which requires kinetic particle-in-cell (PIC) simulations. Semi-analytic modeling of the plasma structure during the discharge phase shows a frequency spectrum $E_{\parallel}(\omega)\approx E_{\parallel,0}(\omega/\omega_{0})^{-1/2}$, where $\omega_{0}$ is an IR scale, taken to be $\omega_{0}\sim 1/h$. Assuming spatial fluctuations have the same spectrum [^5] See [Caputo:2023cpv] for a plausibility argument for this assumption and a statement of the requirement of simulations for definitive justification.[Caputo:2023cpv] demonstrated that in the small-$k$ limit, relevant for gravitationally bound axions, the effect of small-scale oscillations leads to a suppression of $\sim(m_{a}h)^{-4}$ in the axion luminosity [^6] The expression quoted in [Caputo:2023cpv] is $(m_{a}h)^{-3}(\omega_{0}/m_{a})$, but we assume $\omega_{0}\sim 1/h$.. Putting together the high-mass and low-mass regimes, we obtain
@@
where $\Theta(x)$ is the Heaviside theta function, which selects axions with velocity less than the escape velocity of the neutron star, $v_{\rm esc}=\sqrt{2GM_{\rm NS}/R_{\rm NS}}\approx 0.5c$. A simplifying assumption made in (2.22) is that the differential axion production rate is approximately flat at small $k$. The distribution of axion velocities near the NS surface was studied using ray-tracing simulations in [Caputo:2023cpv], and is approximately flat at low-$k$, before dropping at about $v\approx 0.9v_{\rm esc}$. Such simulations are required to make more robust predictions about axion cloud formation and dissipation. The total luminosity of axions emitted into bound states is then
@@
Equation (2.23) agrees with the result in [Caputo:2023cpv] up to an order unity factor that arises from the approximation $d^{3}k\sim 4\pi\bar{k}^{3}/3\sim 4\pi m_{a}^{3}/81$, where $\bar{k}\sim m_{a}/3$ is the average cloud axion momentum.
@@
The axion cloud signal appears as a broadened spectral line centered at frequency corresponding to the axion mass, with natural bandwidth set by the axion velocity distribution (see Eq. (2.26)). Several effects can broaden this signal, similar to the case of radio emission from axion DM, but these are generally subdominant to the natural width. For instance, asphericity of the conversion surface produces both Doppler and gravitational broadening. The Doppler effect arises from co-rotation of the magnetosphere: at different radii, the plasma’s linear velocity projected along the line of sight varies. The typical width scales as $(\Delta f/f)_{\rm Doppler}\sim\Omega,\Delta r$, where $\Delta r$ is the difference between the minimum and maximum radii of the conversion surface. Since conversion generally does not occur for $r>10,R{\rm NS}$, the maximal Doppler contribution is $(\Delta f/f)_{\rm Doppler}\sim 2\times 10^{-3}$ for a pulsar with rotation period $P=1$ s—well below the natural width. Gravitational redshift adds an additional broadening of order $(\Delta f/f)_{\rm grav}\sim\sqrt{1-2GM_{\rm NS}/r_{\rm max}}-\sqrt{1-2GM_{\rm NS}/r_{\rm min}}$. This is typically at the sub-percent level for order-unity deviations from sphericity, but can approach $\sim 10\%$ for highly aspherical surfaces that extend very close to the neutron star surface. A more detailed treatment of these effects is left to future work; here we proceed with the estimate in (2.26). For generality, the scaling equation above has assumed a constant signal bandwidth $\mathcal{B}\sim 60\,\text{MHz}$, suitable for both a population survey and a fixed target observation; the latter case could be made more precise by incorporating the full parametric dependence as given in Eq. (2.26). Note that in the DSA-2000 frequency range $0.7\,\mathrm{GHz}\lesssim f\lesssim 2\,\mathrm{GHz}$ (corresponding axion masses $2.9\,\mu\mathrm{eV}\lesssim m_{a}\lesssim 8.3\,\mu\mathrm{eV}$), the axion wavelength is almost always less than the gap height $h$ estimated by Eq. (2.18), such that taking $\min{(1,{(m_{a}h)}^{-4})}\approx{(m_{a}h)}^{-4}$ is justified in the estimate of $\dot{E}_{a}$  [^7] In this regime, an additional radio line is sourced due to the oscillating current associated with the axion field, which may have interesting observational consequences [Caputo:2023cpv]. The magnitude of the signal is primarily determined by the axion-photon coupling strength, the NS surface magnetic field, its distance to Earth, and its rotation period. Optimal candidates are nearby, fast-spinning pulsars with strong magnetic fields.
@@
The preceding section focuses on radio signals from individual NSs. Another approach to axion detection involves searching for signals from populations of NSs [Safdi:2018oeu, Foster:2022fxn, Bhura:2024jjt]. An advantage of population signals is that the total received flux is bounded from below by the flux from a single NS. However, in the case of axion DM, the signal is Doppler broadened to $\mathcal{B}/f\sim v_{\rm gal}/c\approx 7\times 10^{-4}$, where $v_{\rm gal}$ is the galactic velocity dispersion. Therefore, the expected signal is much broader than that of an individual NS’s $\mathcal{B}/f\sim v_{0}\approx 10^{-6}$ [^8] Even signals from individual NSs are expected to have bandwidth greater than $\sim 10^{-6}$ due to the NS’s rotation (see, e.g., [Witte:2021arp]), but still less than the Doppler-broadened population signal.. Here, we consider radio signals from a population of NSs surrounded by axion clouds. The expected bandwidth for axion cloud signals in (2.26) dominates over Doppler broadening. Therefore, we can estimate the total flux density received from a population of NSs by simply summing the luminosity density of each individual NS:
@@
Figure 6: Distribution plots of the population model (dots), the ATNF catalog (triangles), the Crab pulsar (circled triangle), and the best predicted pulsar to be discovered by the DSA-2000 (circled dot). The color bar shows the axion cloud signal flux density from each individual source, normalized to the flux density from the Crab pulsar. The grey lines show the contours of the axion cloud signal flux density in the corresponding parameter space, normalized to the axion cloud signal flux density from the Crab pulsar. The contour lines are calculated with the varying parameters shown in the plots ($B_{0},d$ for the left panel and $B_{0},P$ for the right panel) and the rest of the parameters fixed to be identical to that of the Crab pulsar. We have taken benchmark values of $m_{a}=5\,\mu\text{eV}$ and $g_{a\gamma\gamma}=10^{-15}\,\text{GeV}^{-1}$ in our estimation of the flux, and the Crab pulsar signal flux density is $F_{\text{Crab}}=1.6\times 10^{-9}\,\text{Jy}$. The left panel shows the distribution in surface magnetic field ($B_{0}$ [G]) and distance to Earth ($d$ [pc]). The distribution plot is bounded from above by the size of the Milky Way, at $d\sim 1.5\times 10^{4}$ pc. The right panel shows the distribution in surface magnetic field ($B_{0}$ [G]) and rotation period ($P$ [s]). Equivalently, one can obtain the distribution on a $P\dot{P}$ diagram using the relation in Eq. (2.5). The gray region is where the conversion radius $r_{c}$ falls inside the NS. We have discarded NSs older than $10^{7}$ yrs in the population model due to their deviations in the magnetosphere from the GJ model. This age cut keeps the population (dots) away from the region with high period $P$ and weak magnetic field $B_{0}$.
@@
For our population modeling, we rely on a pulsar population snapshot generated from the PrsPopPy simulation [Bates:2013uma], which provides NS parameters, including their Period $P$ and distance $d$, and only includes objects discoverable with DSA-2000. We generate their present-day surface magnetic field, $B_{0}\,$[G], by drawing from a log-normal distribution centered at $11.98$ and with a standard deviation of $0.52$. These parameters are consistent with the PsrPopPy simulation [Bates:2013uma], as well as the ATNF catalog [Manchester:2004bp]. Since the magnetosphere of old pulsars is not well-described by the GJ model [Electrosphere], we discard the NSs with age $t_{\text{NS}}\geq 10^{7}\,\text{yrs}$. We also do not include millisecond pulsars with rotation periods $P\lesssim 0.05\,\text{s}$. After imposing the above additional criteria on the population model for new detections, we are left with $\approx 9000$ pulsars that satisfy those criteria, which we use for the axion cloud forecast. In Fig. 6, we show the distributions of the expected detection population (dot) along with the ATNF catalog (triangle). The axion cloud signal flux density from these sources, normalized to the Crab pulsar (circled triangle), is shown in the color bar. The left panel shows the distribution in surface magnetic field, $B_{0}$ and distance $d$, where we have cut off NSs outside of the Milky Way ($d\gtrsim 1.5\times 10^{4}\,$pc). The right panel shows the distribution in surface magnetic field and pulse period PP [^9] Equivalently, one can obtain the distribution on a $P\dot{P}$ diagram using the relation in Eq. (2.5). Here we show a $PB_{0}$ diagram instead to illustrate the effect of surface magnetic field strength, a dominant factor affecting the signal flux density., where we cut off the region where the conversion surface falls inside the NS surface ($r_{c}<R_{\mathrm{NS}}$), as well as NSs older than $10^{7}\,$yrs. We also plot contours of the signal flux density $F_{\mathrm{NS}}$ in the corresponding parameter space with benchmark parameters $m_{a}=5\,\mu\mathrm{eV}$ and $g_{a\gamma\gamma}=10^{-15}\,\mathrm{GeV}^{-1}$ (while fixing the other parameters to be identical to the Crab), normalized against the Crab pulsar ($F_{\mathrm{Crab}}\approx 1.6\times 10^{-9}\,\mathrm{Jy}$). We have chosen the above described cuts on distance, period, and magnetic field for the NSs included in our forecast to be conservative. However, once DSA-2000 discovers new pulsars, the assumptions on the new population made here can be updated and validated before using the collected data to set a constraint.
@@
We also show the best predicted target from the population model (circled dot, parameters shown in Table 3), which represents a prospective candidate that could be discovered by the DSA-2000, whose axion cloud signal flux density is higher than that of the Crab pulsar but has not been observed before. The best predicted target from the population model has a stronger magnetic field, which enables a stronger axion cloud signal than the Crab pulsar. However, due to its larger distance from the Earth, the new candidate produces a smaller non-axion radio flux density than the Crab pulsar [Tolman_2022]. We expect data from the DSA-2000 to discover targets farther from the Earth that could be potential better candidates for observing axion signals. We note that the parameters shown in Table 3 for the best predicted target, informed by our population modeling, are speculative. The data to be taken with the DSA-2000 will ultimately unveil the best discovered target, whose measured parameters will determine the sensitivity to the QCD axion band that can be reached with targeted follow-up observations.
@@
Figure 7: Projected QCD axion discovery reach (SNR $\geq 5$) with the DSA-2000 from radio emission of axion clouds around NSs. We make projections of the Crab pulsar (dashed magenta, $\Delta t_{\text{obs}}=10$ hours), the ATNF population (solid green), and the population model forecasting the DSA-2000 detection (solid blue), based on the planned all sky survey which will observe each NS for 4 hours. We also include projections of a predicted candidate that could potentially be discovered by the DSA-2000 (see Sec. 2.4 for more discussions), with a $\Delta t_{\text{obs}}=10$ hours (dashed red) and a $\Delta t_{\text{obs}}=1000$ hours (dotted red) observation. Other relevant astrophysical constrains are plotted in darker gray (dominated by unbounded polar cap axions from NS [Noordhuis:2022ljw]), and relevant haloscope constraints (ADMX [Asztalos_2010, ADMX:2018gho, ADMX:2018ogs, ADMX:2019uok, ADMX:2021mio, ADMX:2021nhd, ADMX:2024xbv], CAPP [Lee:2020cfj, Jeong:2020cwz, CAPP:2020utb, Lee:2022mnc, Yoon:2022gzp, Kim:2022hmg, Yi:2022fmn, Yang:2023yry, Kim:2023vpo, CAPP:2024dtx], RBF+UF [PhysRevLett.59.839, Wuensch:1989sa, PhysRevD.42.1297, Hagmann:1996qd]) on axion DM are plotted in lighter gray. The QCD axion parameter space is shaded in yellow [Saikawa:2024bta].